The ratio of the length and breadth of a rectangular farm is 2:1 . If its area is 450 m^2, find its length.
Answer by ad_alta(240)

Let 'l' be the length and 'w' the width. Then l/w=2 and lw=450. Substituting w=l/2 into the second equation, we get l(l/2)=450. Therefore l=30.